# An Empirical Study of Developer-Provided Context for AI Coding Assistants in Open-Source Projects

### v1

@@ Abstract.
This paper presents a large-scale empirical study to characterize this emerging form of developer-provided context. Through a qualitative analysis of 401 open-source repositories containing cursor rules, we developed a comprehensive taxonomy of project context that developers consider essential, organized into five high-level themes: Conventions, Guidelines, Project Information, LLM Directives, and Examples. Our study also explores how this context varies across different project types and programming languages, offering implications for the next generation of context-aware AI developer tools.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/06copyright: noneccs: Software and its engineering Software development techniquesccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
While early research has focused on helping developers more effectively prompt LLMs for code generation (nam2025promptingllmscodeediting; yang2025prompts; pickering2025howhumans), more recently, developer-authored project-specific directives for AI agents have emerged as an effective approach (cursor2025rules; windsurf2025rules; copilot2025rules). These “rule files,” exemplified by the cursor.md files in the Cursor IDE (See Figure 1), are distinct from ephemeral, single-shot prompts; they are authored by the end-user for a machine collaborator, with the purpose of providing global context.
@@
Figure 2. Taxonomy of context types in cursor rules.
This paper presents a large-scale empirical study to characterize this new form of developer-provided context. We conducted a qualitative study of a curated dataset of 401 open-source repositories containing developer-authored rule files, seeking to understand what aspects developers consider essential for AI first. And then, we explored how this context varies across different project types and programming languages.
Our analysis reveals that developers provide a rich variety of context types, which we organized into five high-level themes: Conventions, Guidelines, Project Information, Examples and LLM Directives. We also found that many of the rules overlap with existing software documentation practices, such as contribution guidelines and project overviews, suggesting that developers are adapting traditional documentation norms to the new context of AI-assisted development. Meanwhile, we observed unique patterns specific to AI directives, such as explicit instructions targeted at LLM behavior, which have no direct analog in human-oriented documentation but more closely resemble prompt engineering techniques.
The contributions of this paper are threefold:
  • We present a large-scale empirical study of developer-authored rule files for AI coding assistants, providing a systematic characterization of the types of context developers consider essential.
  • We developed a comprehensive taxonomy of context types (Figure 2), organized into five high-level themes, which can inform the design of future context-aware AI developer tools.
  • We explored how context provision varies across different project types and programming languages, offering insights into the different needs of software developers when working with AI assistants.
@@ 2.1. Context Engineering
In recent years, with the advancement of LLMs, context engineering has become a critical area of research. Many studies have shown that the performance of LLMs is highly dependent on how tasks are communicated, i.e., prompt engineering, and the context provided to the model (feldman2024ragged; kojima2022large). Significant attention has been given to different prompting techniques, such as chain-of-thought prompting (wei2022chain; kojima2023large), self-consistency (wang2023self), zero-shot (kojima2022large), and few-shot learning (brown2020language). These techniques demonstrate that providing examples or instructions in the prompt can significantly improve model performance on various tasks. Retrieval-augmented generation (lewis2020retrieval; guu2020retrieval; shuster2021retrieval) also has been widely adopted to enhance context by incorporating relevant information from external sources, such as documents or databases, into the model’s input. This approach allows LLMs to access a broader range of information and improve their performance on specific tasks. More recently, efforts have focused on creating more structured and persistent forms of context that can be easily retrieved and utilized by the model during interactions. One promising direction is the development of memory-augmented models, which can maintain a long-term memory of past interactions and user preferences (shaikh2024aligning; zhang2018personalizing) to provide personalized and context-aware responses.
@@
A similar idea has also been explored in the software engineering domain, where researchers have investigated various techniques to augment the context provided to LLMs for code generation and understanding tasks. For example, Mathews et al. (mathews2024test) found that including test cases enhances function-level code generation in benchmarks such as MBPP (austin2021program) and HumanEval (chen2021evaluating). Jiang et al. (jiang2024self) showed that dividing the coding process into planning and implementation phases provides valuable guidance for large language models. Similar ideas were demonstrated to be useful for code summarization (ahmed2024automatic), repository-level code completion (li2024enhancing), and program repair (autorecover). RAG has also been applied in software engineering tasks, where relevant code snippets, documentation, and abstraction information are retrieved to supplement the model’s context (zhong2025rag; zhang2025coderag; liu2024graphcoder; rag6). This has allowed LLMs to access a broader range of information and improve their performance on code-related tasks. However, most approaches have focused on enhancing LLM performance in tasks that are well-defined and self-contained, such as code completion or code summarization, as covered by existing benchmarks.
Research in IDE settings, where many real-world users interact with LLMs for software development tasks, is still limited. Most efforts are driven by industry players, such as GitHub Copilot (copilot2025rules) and Cursor (cursor2025rules). Although not formally studied, these tools have introduced various techniques to provide context to LLMs in IDE settings, such as retrieving open files, linter errors, or allowing developers to specify files or code snippets. needmorecontext conducted a survey of 131 software professionals using LLMs for coding tasks and found that more context is necessary, but did not identify what specific context is needed. Similarly, contextnotpersonalized found that contextual information is helpful, but LLMs may not know where to focus.
The purpose of providing context to LLMs should not focus solely on fixing coding errors or improving coding accuracy. Other aspects—such as workflow, usability, maintainability, communication, and alignment with project goals—are equally important in software development, yet have not been well studied in prior work. developerneed1 conducted a large-scale difference-in-differences study of Cursor’s impact on software projects and found that projects adopting Cursor suffered from an increase in static analysis warnings and code complexity, indicating higher technical debt in the future. Several studies have assessed the usability issues of AI coding assistants (developerneed2; developerneed3), finding that developers often struggle to effectively use AI coding assistants for certain tasks, such as generating test cases or writing natural language artifacts in interview settings. Other studies have attempted to build fully automated AI coding assistants with predetermined workflows (fullyautomate3; fullyautomate1; fullyautomate2), but these typically do not consider that workflows may vary significantly across disciplines or projects. Multi-agent systems have also been proposed to handle complex software engineering tasks; nevertheless, these approaches often emphasize the need for more explicit context from developers to guide the agents (agentneedcontext), or rely on fabricated context (lackcontext). However, previous studies may have failed to recognize what developers actually need or want to convey to AI coding assistants in real-world settings. Our study is the first to systematically investigate the context developers consider essential when working with AI coding assistants in real-world software projects, using cursor rule files as a proxy.
@@ 2.2. Rules for AI Coding Assistants
Many AI coding assistants have introduced ways for developers to author system-level instructions that provide globally available context, preferences, or workflows for the underlying AI agents to follow. These directives, often referred to as “rule files,” can include project-specific, team-level, or user-specific guidelines that guide the behavior of AI coding assistants. For example, Cursor, an AI-powered code editor, allows developers to create “cursor rules” in ‘.mdc‘ files that specify guidelines and constraints for code generation (cursor2025rules). Because LLMs, by default, do not retain memory between sessions and may not know where to focus when generating code, these rules provide a way to persistently encode important context that the LLM should consider. When a developer includes these rules in their project, AI coding assistants can incorporate them as part of the context when generating responses, leading to more relevant and accurate suggestions.
@@ 2.3. Documentation and Communication in Software Engineering
Software engineering is an information-intensive and collaborative activity, making effective documentation and communication critical to the success of software projects. Numerous studies have investigated various aspects of software documentation, including the types of information developers seek (brandt2009two; maalej2013patterns; freund2015contextualizing; meng2018application), as well as tools and techniques to improve documentation practices (robillard2011field; aghajani2019software). For example, Maalej and Robillard (maalej2013patterns) conducted a large-scale study of API documentation and identified common patterns of information needs among developers, finding twelve distinct types of information frequently sought in API documentation. Similarly, other studies have explored the types of knowledge used in GitHub README files (prana2019categorizing; fronchetti2023contributing), categorizing the information developers use when collaborating on open-source projects. These studies have provided valuable insights into the information needs of developers and the challenges they face in accessing and utilizing documentation effectively. While we have a good understanding of the types of information developers need and provide when working with human collaborators, less is known about how they communicate with AI agents. Since AI coding agents have only recently become widely adopted and started to be considered collaborators, the ways developers document their needs for these agents remain largely unexplored. In this paper, we build upon prior work by focusing specifically on the context provided by developers for AI coding assistants, and how this context differs from traditional documentation practices.
@@ 3. Methodology
Figure 3. Overview of the qualitative coding process used to develop the taxonomy of cursor rule codes and the coded rules for quantitative analysis.
To understand the types of context developers provide when authoring cursor rules, we conducted a qualitative study of cursor rules from real-world open-source software repositories. A brief overview of our coding process is illustrated in Figure 3. We selected cursor rules because they were among the earliest to introduce this feature, have been widely adopted in the open-source community, and are frequently discussed in developer forums such as Reddit  [^1] https://www.reddit.com/r/cursor/comments/1ju63ig/the_one_golden_cursor_rule_that_improved/ and Hacker News  [^2] https://news.ycombinator.com/item?id=43305919.
@@ 3.1. Data Collection and Preprocessing
Data Fetching. For data fetching, we used Sourcegraph [^3] https://sourcegraph.com/ client. Sourcegraph is an open-source code search tool that indexes all public GitHub repositories with at least one star, providing a good representation of the most popular open-source repositories on GitHub. We fetched repositories containing at least one cursor rule file (i.e., with the .mdc extension), excluding any forked or archived repositories, which resulted in 487 repositories. We cloned all the repositories locally and fetched metadata using GitHub’s official API for further analysis.
Data Cleaning. We removed 15 repositories that did not contain any .mdc files after cloning (e.g., the files may have been deleted in later commits). We then manually checked the content of the remaining repositories and removed 11 repositories whose .mdc files were used for other purposes, such as SD card configuration. Next, we manually checked whether the .mdc files in the remaining repositories were written in English and removed 43 repositories that were not. This left us with 418 repositories for further analysis.
Since some repositories contained only empty data or were exact duplicates of others, we removed 12 such repositories after manual review. Additionally, some repositories included only template or example .mdc files intended for others to use, such as awesome-cursor-rules-mdc [^4] https://github.com/PatrickJS/awesome-cursorrules. Because the main purpose of our study is to analyze the usage of cursor rules in real-world applications, we removed five such repositories after manual review, leaving 401 repositories for further analysis.
@@ 3.2. Dataset
Among all repositories, we identified a total of 1,876 .mdc files. On average, each repository contained 4.68 .mdc files (SD = 6.89), with the maximum number of .mdc files in a single repository being 56. According to the latest cursor rule documentation, it is recommended to place all cursor rule files in a dedicated directory at the root of the repository, called .cursor. We found that 374 repositories (93.27%) followed this recommendation. In these repositories, an average of 1.72 authors contributed to the directory (SD = 1.32), with a maximum of 8 authors. The average number of commits to the cursor rule directory was 5.68 (SD = 9.31), with a maximum of 98 commits. Among the 281 repositories that had at least two commits in the cursor rule directory, people spent an average of 16 days (SD = 15.90) between each commit updating their cursor rules. The average length of all cursor rule files was 462.67 lines (SD = 1,197.00), with the longest file containing 11,076 lines.
The entire dataset, prompt, reproduce package is available via this link [^5] https://anonymous.4open.science/r/cursorrule-supp-110E.
@@ 3.3. Qualitative Analysis
We qualitatively analyzed the content of cursor rules to understand the types of context developers provide when authoring these files. We employed thematic analysis, following established procedures (braun2006using; clarke2013teaching), to systematically identify recurring patterns and themes. The analysis was conducted iteratively, involving multiple rounds of coding, discussion, and refinement among the research team.
We specifically focused on analyzing the content of cursor rules, excluding code blocks and other non-textual elements, such as images, since the purpose and content of these non-textual elements were mostly self-explanatory. For example, code blocks were typically included to provide code examples or snippets, which did not require further interpretation. We programmatically excluded these non-textual elements from our analysis to focus on the textual content.
We, the two authors, began by familiarizing ourselves with the data, reading through cursor rules in the selected repositories. Given that analyzing each cursor rule file in all 401 repositories would be impractical, we selected 30 repositories from the entire pool of 401 repositories for open coding. To ensure diversity, we sampled repositories across distinct domains and programming languages. We fetched the list of topics each repository owner assigned to their repository using GitHub’s official API. We then selected the top 10 most popular topics, which are ”AI”, ”TypeScript”, ”tag-production”, ”Java”, ”security”, ”Prisma”, ”ecommerce”, ”MCP-Server”, ”PHP”, and ”DeepSeek”. To prevent class imbalance, we randomly selected 3 repositories from each topic, resulting in a total of 30 repositories. We carefully examined the cursor rules, their descriptions, and related repository documentation to understand the context in which they were written. This process typically required several hours per repository. During this initial review, we decided to code at the line level, as developers often expressed a single idea per line in cursor rule files. As there were lines that were not directly related to context provision, such as section headers or empty lines, we introduced ”No Code” to indicate they did not contain relevant context information. Guided by these insights, we generated initial codes to describe the purpose and content of each cursor rule line through an inductive approach. We recorded notes on emerging impressions and potential themes, referencing prior literature on documentation, developer communication, and context engineering to inform our perspective.
@@ 3.4. Threats to Validity
A further limitation arises from the use of a large language model (LLM) to assist in the coding process. To improve reliability, we repeated the labeling process three times, selected the majority vote as the final label, and found substantial agreement with human coding on the initial 30 repositories. Nevertheless, there remains a possibility that the LLM introduced biases or misinterpretations, and its outputs may not fully capture the nuances, context, or intent behind developers’ cursor rules. Such potential bias is a common limitation of qualitative analysis, especially when conducted at scale. Based on a human manual verification, LLM usually disagree with human coders on functionality and language/framework specific rules. If all three labels were different, we marked it as “No Code” to indicate uncertainty. There are 2.1% of all lines fall into this disagreement. However, considering the size of our dataset, this number should be treated as a lower bound.
Our dataset is limited to open-source repositories indexed by Sourcegraph, which focuses on popular GitHub projects. While not exhaustive, this provides a reasonable sample of widely used repositories. In addition, although we focused on cursor rules, other rule formats exist, such as Copilot custom instructions, which may include different types of context. However, since cursor rules were among the earliest to introduce this feature, have been widely adopted in the open-source community, and the mechanisms of other rules are similar to cursor rules, we believe our selection is sufficiently representative to provide initial insights into this emerging practice. Nevertheless, future work could study rules authored for other AI coding assistants to validate and extend our findings.
As part of the data cleaning steps, if multiple repositories contained identical cursor rules, we kept only one and excluded the duplicates (11 repositories) to ensure a comprehensive capture of all unique rules. This approach may underrepresent some rules that are widely copied across different repositories, and it may also influence our subsequent quantitative analysis of rule reuse frequency. However, since only a small portion of repositories are removed due to duplication and most are copied from similar domains, we believe this threat is minimal.

### v2

# Beyond the Prompt: An Empirical Study of Cursor Rules

@@ Abstract.
This paper presents the first large-scale empirical study to characterize this emerging form of developer-provided context. Through a qualitative analysis of 401 open-source repositories with cursor rules, we developed a comprehensive taxonomy of project context that developers consider essential, organized into five high-level themes: Conventions, Guidelines, Project Information, LLM Directives, and Examples. Our study also explores how this context varies across different project types and programming languages, offering implications for the future context-aware AI developer tools.
journalyear: 2026copyright: ccconference: 23rd International Conference on Mining Software Repositories; April 13–14, 2026; Rio de Janeiro, Brazilbooktitle: 23rd International Conference on Mining Software Repositories (MSR ’26), April 13–14, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793302.3793367isbn: 979-8-4007-2474-9/2026/04
@@ 1. Introduction
Early research focused on helping developers prompt LLMs effectively for code generation (nam2025promptingllmscodeediting; yang2025prompts; pickering2025howhumans). More recently, developer-authored, project-specific directives for AI agents have emerged as another approach (cursor2025rules; windsurf2025rules; copilot2025rules). These “rule files,” exemplified by the cursor.md files for the Cursor IDE (see Figure 1), are distinct from ephemeral, single-shot prompts; they are authored by end users for a machine collaborator to provide contextual guidance.
@@
This paper presents the first large-scale empirical study to characterize this developer-provided context. We conducted a qualitative study of 401 open-source repositories containing developer-authored rule files to understand aspects developers consider essential when collaborating with AI. Then, we explored how this context varies across different project types and programming languages.
Our analysis reveals that developers provide a rich variety of context types, which we organized into five high-level themes: Convention, Guideline, Project, Example, and LLM Directive; and 20 detailed codes (Section 4). Many rules overlap with traditional software documentation practices, such as guidelines and project overviews, suggesting that developers are adapting established norms to AI-assisted development. Others, however, reflect patterns unique to AI directives, such as instructions targeting LLM behavior, which have no direct analog in human-oriented documentation and instead resemble prompt engineering techniques. The context provision also varies across different project types and programming languages, reflecting the diverse needs of developers when working with AI assistants. Our work can inform the design of future context-aware AI developer tools and help people better utilize AI coding assistants.
@@ 2.1. Context Engineering
In recent years, with the advancement of LLMs, context engineering has become a critical area of research. Many studies have shown that the performance of LLMs is highly dependent on how tasks are communicated, i.e., prompt engineering, and the context provided to the model (feldman2024ragged; kojima2022large; jiang2025questions). Significant attention has been given to different prompting techniques, such as chain-of-thought prompting (wei2022chain; kojima2023large), self-consistency (wang2023self), zero-shot (kojima2022large), and few-shot learning (brown2020language). These techniques demonstrate that providing examples or instructions in the prompt can significantly improve model performance on various tasks. Retrieval-augmented generation (lewis2020retrieval; guu2020retrieval; shuster2021retrieval) also has been widely adopted to enhance context by incorporating relevant information from external sources, such as documents or databases, into the model’s input. This approach allows LLMs to access a broader range of information and improve their performance on specific tasks. More recently, efforts have focused on creating more structured and persistent forms of context that can be easily utilized by the model during interactions. One promising direction is the development of memory-augmented models, which can maintain a long-term memory of past interactions and user preferences (shaikh2024aligning; zhang2018personalizing) to provide context-aware responses.
@@
A similar idea has also been explored in the software engineering domain, where researchers have investigated various techniques to augment the context provided to LLMs for code generation and understanding tasks. For example, Mathews et al. (mathews2024test) found that including test cases enhances function-level code generation. Jiang et al. (jiang2024self) showed that dividing the coding process into planning and implementation phases provides valuable guidance for LLMs. Similar ideas were demonstrated to be useful for code summarization (ahmed2024automatic), code understanding (nam2024using), repository-level code completion (li2024enhancing), and program repair (autorecover). RAG has also been applied in software engineering tasks, where relevant code snippets, documentation, and abstraction information are retrieved to supplement the model’s context (zhong2025rag; zhang2025coderag; liu2024graphcoder; rag6). This has allowed LLMs to access a broader range of information and improve their performance on code-related tasks. However, most approaches have focused on enhancing LLM performance in tasks that are well-defined and self-contained, such as code completion or code summarization, as covered by existing benchmarks.
@@ 2.2. Rules for AI Coding Assistants
Many AI coding assistants have introduced ways for developers to author system-level instructions that provide globally available context, preferences, or workflows for the AI to follow. These directives, often referred to as “rule files,” can include project-, team-, or user-specific guidelines that shape AI behavior. For example, Cursor, an AI-powered code editor, allows developers to create “cursor rules” in ‘.mdc’ files specifying guidelines and constraints for code generation (cursor2025rules). Because LLMs, by default, do not retain memory between sessions and may not know where to focus, these rules persistently encode important context, enabling AI assistants to incorporate them and generate more relevant and accurate suggestions.
@@
Figure 2. Overview of the qualitative coding process for developing the taxonomy and coding rules for quantitative analysis.
@@ 2.3. Documentation and Communication in Software Engineering
Software engineering is an information-intensive and collaborative activity, making effective documentation and communication critical to the success of software projects. Numerous studies have investigated various aspects of software documentation, including the types of information developers seek (brandt2009two; maalej2013patterns; freund2015contextualizing; meng2018application; nam2024understanding), as well as tools and techniques to improve documentation practices (robillard2011field; aghajani2019software; autocomplete2024jiang). For example, Maalej and Robillard (maalej2013patterns) conducted a large-scale study of API documentation and identified common patterns of information needs among developers, finding twelve distinct types of information frequently sought in API documentation. Similarly, studies have explored the types of knowledge used in README files (prana2019categorizing; fronchetti2023contributing), categorizing the information developers use when collaborating on open-source projects. These studies have provided valuable insights into the information needs of developers and the challenges they face in using documentation effectively. While we have a good understanding of the types of information developers need and provide when working with human collaborators, less is known about how they communicate with AI agents. Since AI coding agents have only recently become widely adopted and started to be considered collaborators, the ways developers document their needs for these agents needs further exploration. In this paper, we build upon prior work by focusing specifically on the context provided by developers for AI coding assistants, and how this context differs from traditional documentation practices.
@@ 3. Methodology
To understand the types of context developers provide when authoring cursor rules, we conducted a qualitative study of cursor rules from real-world open-source software repositories. An overview of our coding process is shown in Figure 2. We selected cursor rules because they were among the earliest to introduce this feature, are popular in the open-source community, and are frequently discussed in developer forums such as Reddit  [^1] https://www.reddit.com/r/cursor/comments/1ju63ig/the_one_golden_cursor_rule_that_improved/ and Hacker News  [^2] https://news.ycombinator.com/item?id=43305919.
@@ 3.1. Data Collection and Preprocessing
Data Fetching. We used Sourcegraph [^3] https://sourcegraph.com/, an open-source code search tool that indexes all public GitHub repositories with at least one star, representing the most popular open-source repositories on GitHub. We fetched all repositories containing a cursor rule file (i.e., with the .mdc extension), excluding any forked or archived repositories, resulting in 487 repositories. We retrieved all repositories and metadata using GitHub’s official API for further analysis.
Data Cleaning. After cloning, we removed 15 repositories that contained no .mdc files. Manual inspection of the rest led to the removal of 11 whose .mdc files served unrelated purposes (e.g., SD card configuration) and 43 whose files were not written in English. We also removed 12 repositories that contained only empty data or exact duplicates. Finally, we excluded five repositories that provided only template or example .mdc files, as our study focuses on real-world usage, leaving 401 repositories for analysis.
@@ 3.2. Dataset
Among all repositories, we identified a total of 1,876 .mdc files. On average, each repository contained 4.68 .mdc files (SD = 6.89), with a maximum of 56 .mdc files in a single repository. The latest cursor rule documentation recomments to place all rule files in the .cursor directory at the root of the repository. Among the 374 repositories (93.27%) followed this recommendation, an average of 1.72 authors contributed to the directory (SD = 1.32), with a maximum of 8 authors. The average number of commits was 5.68 (SD = 9.31), with a maximum of 98 commits. Among the 281 repositories containing the cursor rule directory that has been modified after creation, people spent an average of 16 days (SD = 15.90) between each commit. The average length of all cursor rule files was 462.67 lines (SD = 1,197.00), with the longest file containing 11,076 lines.
@@ 3.3. Qualitative Analysis
We qualitatively analyzed the content of cursor rules to understand the types of context developers provide when authoring them. We employed thematic analysis, following established procedures (braun2006using; clarke2013teaching), to systematically identify recurring patterns and themes. The analysis was conducted iteratively, involving multiple rounds of coding, discussion, and refinement by the research team.
We focused our analysis on the textual content of cursor rules, excluding code blocks and other non-textual elements (e.g., images), as their purpose and content were generally self-explanatory. For instance, code blocks were typically included to provide illustrative code examples or snippets and therefore did not require further interpretation. Accordingly, we programmatically removed these non-textual elements to concentrate on the textual content.
We, the two authors, began by familiarizing ourselves with the data, reading through all cursor rules. Given that analyzing all rules from 401 repositories was impractical, we used stratified sampling to select 30 repositories from the entire pool for open coding. To ensure diversity, we sampled repositories across distinct topics assigned by each repository owner. We selected the top 10 most popular topics, which are ”AI”, ”TypeScript”, ”tag-production”, ”Java”, ”security”, ”Prisma”, ”ecommerce”, ”MCP-Server”, ”PHP”, and ”DeepSeek”. To prevent class imbalance, we randomly selected three repositories from each topic. We carefully checked the cursor rules, their descriptions, and related repository documentation to understand the context. During this initial review, we decided to code at the line level, as developers often expressed a single idea per line in cursor rule files. Since some lines were not directly related to context provision, such as section headers, we introduced ”No Code” to indicate those lines. Through an inductive approach, we generated initial codes to describe the purpose and content of each cursor rule line. We recorded notes on emerging impressions and potential themes, referencing prior literature on documentation, developer communication, and context engineering to inform our perspective. Each researcher spent around 40 hours on this process.
@@ 3.4. Threats to Validity
A further limitation arises from using a large language model (LLM) to assist in coding. To improve reliability, we repeated the labeling process three times, selected the majority vote as the final label, and found substantial agreement with human coding. Nevertheless, the LLM may have introduced biases or misinterpretations, and failing to capture the nuances, context, or intent behind developers’ cursor rules. Such potential bias is a common limitation of qualitative analysis, especially when conducted at scale.
Our dataset is limited to open-source repositories indexed by Sourcegraph, which focuses on popular GitHub projects. While this coverage is not exhaustive, it provides a reasonable sample of widely used repositories. Other rule formats, such as Copilot custom instructions, may include different types of contextual information. Since cursor rules were among the earliest to introduce this feature, have been widely adopted in the open-source community, and share similar underlying mechanisms with other rule formats, we believe our selection is sufficiently representative to provide initial insights into this emerging practice. Future work could study rules authored for other AI coding assistants to validate and extend our findings.